# FoleySpace: Vision-Aligned Binaural Spatial Audio Generation

### v1

@@ V-A Dataset
The training dataset for the diffusion models construted in Section IV. We carefully selected 9,212 single-source audio samples from the VGGSound [28] dataset and manually annotated their source locations in a grid format, as shown in Fig. 4, to construct the evaluation dataset VGGSound-Solo  [^1] VGGSound-Solo is provided in the supplementary material, following the original format of the VGGSound dataset.. The training and evaluation sets do not overlap.
@@ V-C2 Quantitative Evaluation
TABLE II: Objective evaluation results on the VGGSound-Solo. * denotes the mono generator in the proposed framework.
@@
TABLE III: The impact of text label on sound source estimation accuracy in inference
@@
TABLE IV: Comparison of binaural audio Quality Based on RIR and HRIR.
@@ V-H Comparison with RIRs-based Method
From the table we can see that for both fine-grained and coarse-grained spatio-temporal schemes, the HRIRs-based approach yields superior audio quality compared to the RIRs-based approach. This demonstrates that the quality of binaural audio is highly dependent on the fidelity of spatial auditory cues: By incorporating the physiological filtering characteristics of the human ear, HRIRs significantly enhance spatial realism and clarity in sound reproduction.

### v2

@@ V-A Dataset
The training dataset for the diffusion models construted in Section IV. We carefully selected 9,212 single-source audio samples from the VGGSound [28] dataset and manually annotated their source locations in a grid format (as shown in Fig. 4) to construct the evaluation dataset VGGSound-Solo [32]. The training and evaluation sets do not overlap.
@@ V-C2 Quantitative Evaluation
TABLE II: Objective evaluation results for VGGSound-Solo. * denotes the mono generator in the proposed framework.
@@
TABLE III: The effect of text labels on sound source estimation accuracy for VGGSound-Solo inference.
@@
TABLE IV: Comparison of binaural audio quality based on RIRs and HRIRs for VGGSound-Solo.
@@ V-H Comparison with RIRs-based Method
From the table, it is evident that for both fine-grained and coarse-grained spatio-temporal schemes, the HRIRs-based approach outperforms the RIRs-based approach in perceived audio quality. This indicates that the quality of binaural audio is highly dependent on the fidelity of spatial auditory cues. By accurately modeling the physiological filtering characteristics of the human ear, HRIRs can achieve a high level of audio realism.